Liquid hexane reacts with gaseous oxygen gas to produce gaseous carbon dioxide and gaseous water . What is the theoretical yield of water formed from the reaction of 64.6 g of hexane and 84.4 g of oxygen gas? Round your answer to 3 significant figures.
Answer #1

2C6H14 + 19O2 ---------> 12CO2 + 14H2O

Moles of Hexane = mass/molar mass = 64.6/86 = 0.7512

Moles of O2 = 84.4/32 = 2.6375

From reaction,

2 moles Hexane requires 19 moles O2

So, 0.7512 moles will require 19*0.7512/2 = 7.1364 moles of O2

Since available moles of O2 < required moles of O2

Hence, O2 is limiting reagent and it will drive the yield of products.

Again from reaction,

19 moles O2 produces 14 moles H2O

So, 2.6375 moles will produce 14*2.6375/19 = 1.9434 moles of H2O

Mass of H2O (theoretical yield) = moles * molar mass = 1.9434*18 = 34.982 grams OR approx 35 grams .. Answer